### v2

# Hunting for extra dimensions in black hole shadows

Abstract
Observational data of the Sagittarius A* (Sgr A*) shadow released by the Event Horizon Telescope are used to investigate eventual deviations in the black hole shadow radius, aiming to seek physics beyond the Standard Model coming from extra-dimensional theory. We consider the braneworld scenario described by the Randall-Sundrum model and determine the black hole shadow radius correction owing to the higher dimension. From data of the shadow radius in units of black hole mass determined by Keck- and Very Large Telescope Interferometer (VLTI)-based estimates, we imposed restrictions on the deviation obtained, and one sets an upper limit to the curvature radius of anti-de Sitter ($\mathrm{AdS_{5}}$) spacetime $\ell_{\left(\text{Keck}\right)}\lesssim 3.8\times 10^{-2}\,\mathrm{AU}$ and $\ell_{\left(\text{VLTI}\right)}\lesssim 1.6\times 10^{-2}\,\mathrm{AU}$ (at $68\%$ confidence level), respectively.
extra dimension, brane, black hole, quasinormal modes
pacs:
04.50.−h, 04.50.Gh, 11.10.Kk
I Introduction
Over the past century, attempts have been made to investigate the geometrical structure of the Universe from the proposal of the existence of extra dimensions (Nordstrom:1914ejq, ). Following the development of general relativity (GR), a theoretical extra-dimensional model was built, which aimed to unify the electromagnetic and gravitational fields (Kaluza:1921tu, ; Klein:1926tv, ). Recently, the modern braneworld theories have renewed interest in higher dimensional theories (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ; Randall:1999ee, ; Randall:1999vf, ). These proposed models arise in the unification schemes, where one expected that the unification of gravitational and gauge forces would occur near the $\mathrm{TeV}$ scale (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ). In such a case, the gauge interactions, and therefore the ordinary matter, get confined to a four-dimensional hypersurface (three-brane), whereas gravity lives in a higher-dimensional spacetime known as bulk (Randall:1999vf, ; Randall:1999ee, ). Thus, gravity would be diluted in the higher-dimensional space, spreading throughout the bulk, explaining, therefore, the observed weakness of gravity compared to other fundamental interactions of nature (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ; Randall:1999ee, ; Randall:1999vf, ). Consequently, this is a promising scenario to address the hierarchy problem that deals with the discrepancy between the gravitational and electroweak fundamental energy scales without relying on low-energy supersymmetry or technicolor (Antoniadis:1998ig, ; Arkani-Hamed:1998jmv, ; Arkani-Hamed:1998sfv, ).
In this framework, although the Standard Model (SM) fields are trapped on our four-dimensional manifold, events of high-energy collisions can, for instance, eventually cause the SM particles to be kicked off into the new dimensions (ParticleDataGroup:2022pth, ). Thus, collider experiments have been employed to seek evidence of large extra dimensions in the Arkani-Hamed-Dimopoulos-Dvali (ADD) model, and it was found that the most stringent limits on the Planck mass of higher-dimensional spacetime come from the ATLAS (CMS) missing energy experiment, giving $M_{D}>5.9–11.2\,\mathrm{TeV}$ (ATLAS:2021kxv, ). In its turn, analysis from astrophysical and cosmological origin, such as light Kaluza-Klein (KK) graviton production in stars, has also provided the strongest bounds so that $M_{D}>1700\,(76)\,\mathrm{TeV}$ for $\delta=2\left(3\right)$ (Hannestad:2003yd, ). Further discussions have attempted to seek traces of extra dimensions in several physical systems, such as in rotating torsion balance (Adelberger:2009zz, ; Murata:2014nra, )—wherein one tests Newton’s law at sub-mm distances—spectroscopy and atomic physics (Dahia:2015xxa, ; Dahia:2015bza, ; Dahia:2017jya, ; Lemos:2019rhj, ; Lemos:2021xpi, ; Luo:2006ad, ; Zhi-gang:2007swh, ; Salumbides:2015qwa, ; Lemos:2022vhw, ), and black hole production (CMS:2018ozv, ; ATLAS:2015yln, ).
On the other hand, the model proposed by Randall and Sundrum (RS) (Randall:1999ee, ; Randall:1999vf, ) describes our Universe by postulating it as a five-dimensional anti-de Sitter ($\mathrm{AdS_{5}}$) spacetime with a warped geometry. In this scenario, bounds on the $\mathrm{AdS_{5}}$ curvature radius $\ell$ have also been imposed from astrophysical observations. For instance, the investigation of the persistence of x-ray binaries containing a black hole in an $\mathrm{AdS_{5}}$ braneworld shows that $\ell\lesssim 10^{-2}\,\mathrm{mm}$ (Emparan:2002jp, ). Furthermore, recent work has also aimed to search for signals of the extra dimension on events of black hole gravitational waves (Bohra:2023vls, ; Chrysostomou:2022evl, ). In such a way, the study of Hawking radiation emitted by stellar-mass black holes found that $\ell\lesssim 5\,\mathrm{\mu m}$ from gravitational wave measurements (McWilliams:2009ym, ), while Ref. (Vagnozzi:2019apd, ) —analyzing the shadow of M87*—imposes constraints on the $\mathrm{AdS_{5}}$ radius such that $\ell\lesssim 170\,\mathrm{AU}$. The multimessenger cosmology, which considers joint observations of gravitational waves and electromagnetic counterparts, has been used to set limits on the radius of curvature of the higher-dimensional spacetime, $\ell\lesssim 0.535\,\mathrm{Mpc}$ (Visinelli:2017bny, ). On its turn, Ref. (Mishra:2021waw, ) constrains the tidal charge of a rotating braneworld black hole, whereas in Ref. (Hou2021, ) one investigates how the extra dimension affects the shape of the black hole shadows.
It is known that the black holes only recently have been experimentally probed. Thus, one expect that black hole physics may provide us with significant insights into the structure of spacetime (Chakraborty:2017qve, ; Dey:2020pth, ; Hashemi:2019jlt, ; Anacleto:2023ntm, ; Cruz:2020emz, ; Cardoso:2003cj, ; Konoplya:2011qq, ; Konoplya:2003ii, ), as well as allow us to test fundamental physics (Yagi:2016jml, ; Goddi:2016qax, ). In this context, recent measurements of the Event Horizon Telescope (EHT) provided images of supermassive black hole shadow at the M87* galaxy center (EventHorizonTelescope:2019dse, ; EventHorizonTelescope:2019ggy, ), in addition to the shadow of compact source Sagittarius A* (Sgr A*), a supermassive black hole in the center of the Milky Way (EventHorizonTelescope:2022xqj, ; EventHorizonTelescope:2022wkp, ). It is noteworthy that, in this scenario, the search for physics beyond the SM has been based on analysis of the shadow radius and deviation from the circularity $\Delta C$ of the black hole shadow (vagnozzi2022horizon, ). However, whereas the limit on the circularity deviation has been quoted for the M87* measurements, imposing $\Delta C\lesssim 10\%$ (EventHorizonTelescope:2019dse, ), such information for Sgr A* is not available. Therefore, the approach employed in Ref. (Vagnozzi:2019apd, ) cannot be applied to test fundamental physics and to put limits on the $\mathrm{AdS_{5}}$ curvature radius from the circularity deviations of Sgr A* (Vagnozzi2022, ). Instead, one must proceed with an investigation based on its shadow size.
Since the extra dimensions have been proposed in modified frameworks of GR, one expects that they might yield signals on the black hole shadow. Therefore, black hole images can be used to test fundamental physics (Yagi:2016jml, ; Vagnozzi2022, ). In this work, we perform a study of the shadow radius of Sgr A* aimed at obtaining new and complementary constraints on the extra-dimensional $\mathrm{AdS_{5}}$ curvature radius from observational data. In this case, we imposed limits on deviations from the GR’s predictions arising from a braneworld theory. For this purpose, we will consider a black hole in the RS background in the weak field regime, described by the Garriga-Tanaka metric (Garriga:1999yh, ). This solution has been obtained by assuming a classic approach that provides the linearized braneworld gravity. As we will see, our analysis allowed us to place the bounds on the extra-dimensional curvature radius $\ell$ in an alternative way, avoiding the issue of lack of constraints on the deviation from circularity of the shadow measures for Sgr A* and exploring its inaccurate spin measurements (Banerjee:2022jog, ; Fragione:2020khu, ).
The organization of this paper is as follows: In Sec. II, by considering the Randall-Sundrum braneworld scenario, we investigate the null geodesic curves aiming to find the shadow radius of the RSII black hole and discuss the instability of the null circular orbits from the Lyapunov exponent. In Sec. III, from the EHT observational data of the Sgr A*’s image black hole, we set the imposed constraints on the $\mathrm{AdS_{5}}$ curvature radius $\ell$. The concluding remarks and the main results are given in Sec. IV.
II Corrected Shadow Radius of the Braneworld Black Hole
In this section, we will study the null geodesic of the Garriga-Tanaka (RSII) black hole, aiming to investigate the influence of the extra dimension on the critical impact and critical radius parameters. This approach will also allow us to determine the radius of the shadow of the braneworld black hole. From this investigation, we intend to find bounds on the curvature radius of the $\mathrm{AdS_{5}}$ spacetime. In this case, we start by considering the Garriga-Tanaka solution that, at the first-order approximation of $GM$, describes the metric of a matter distribution with mass $M$ localized in the brane (Garriga:1999yh, ),
$ds^{2}=-f(r)dt^{2}+\dfrac{dr^{2}}{g(r)}+r^{2}d\Omega^{2},$ (1)
where, in the weak field limit $\left(r\gg\ell\right)$ from the horizon of matter after the collapse, we get
$f(r)=1-\dfrac{2M}{r}-\dfrac{4M\ell^{2}}{3r^{3}},\quad\text{and}\quad g(r)=1-\dfrac{2M}{r}-\dfrac{2M\ell^{2}}{3r^{3}},$ (2)
in which $\ell$ is the curvature radius of $\mathrm{AdS_{5}}$, and we consider unities $c=G=1$. We can find the geodesics of the metric (1), considering the following Lagrangian density
$\mathcal{L}\equiv\frac{1}{2}g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu},$ (3)
where the dot denotes the derivative with respect to the affine parameter. Now, applying the line element (1), we have
$2\mathcal{L}=f(r)\dot{t}^{2}-\dfrac{\dot{r}^{2}}{g(r)}-r^{2}\left(\dot{\theta}^{2}+\sin^{2}\theta\dot{\phi}^{2}\right).$ (4)
For the sake of simplicity, at this point, by setting the angle $\theta=\pi/2$, one can examine the motion in an equatorial plane. Thus, we can find the equation of motion for the Hamilton-Jacobi equation
$E=f(r)\dot{t},\qquad\quad L=r^{2}\dot{\phi},$ (5)
where $E$ and $L$ are the constants corresponding to energy and angular momentum, respectively. For a null geodesic, we have $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=0$, and we can find the following expression
$\dot{r}^{2}+\mathcal{V}(r)=0,\qquad\mathcal{V}(r)=\dfrac{g(r)}{f(r)}E^{2}-g(r)\dfrac{L^{2}}{r^{2}}.$ (6)
Now, just for convenience, we can introduce a new variable $u=1/r$ to write the orbital equation
$\left(\dfrac{du}{d\phi}\right)^{2}=\dfrac{\left(3-6Mu-2M\ell^{2}u^{3}\right)}{\left(3-6Mu-4M\ell^{2}u^{3}\right)b^{2}}-\left(1-2Mu-\dfrac{2M\ell^{2}u^{3}}{3}\right)u^{2},$ (7)
$\dfrac{d^{2}u}{d\phi^{2}}=\dfrac{2M\ell^{2}\left(3-4Mu\right)u^{2}}{b^{2}\left(6Mu+4M\ell^{2}u^{3}-3\right)^{2}}-\left(1-3Mu-\dfrac{5M\ell^{2}u^{3}}{3}\right)u.$ (8)
Here, we have that $b=L/E$ is the impact parameter defined as the perpendicular distance (measured from infinity) between the geodesic and the parallel line that passes through the origin. Using the following critical conditions, (i) $\mathcal{V}(r_{c})=0$ and (ii) $d\mathcal{V}(r)/dr\big{|}_{r=r_{c}}=0$, we can obtain the critical impact parameter from a system of equations with which one relates the impact parameter to the critical radius, as follows:
$b_{c}^{2}=\dfrac{r_{c}^{2}}{f(r_{c})}=\dfrac{3r_{c}^{5}}{3r_{c}^{3}-6Mr_{c}^{2}-4M\ell^{2}}.$ (9)
Notice that $r_{c}$ corresponds to the radius of photon circular orbit. On its turn, the critical conditions expressed by (ii) yield a polynomial equation that is used to determine the critical radius:
$0\displaystyle 20\ell^{4}M^{2}+6\ell^{2}M\left(13M-6r_{c}\right)r_{c}^{2}+9r_{c}^{4}\left(6M^{2}-5Mr_{c}+r_{c}^{2}\right)=$ 0 . (10)
We must highlight that, here, we want to study the trajectory of a ray of light in the spherically symmetric Garriga-Tanaka background. Specifically, when we analyze the ray in a plane, we observe that any ray of light that starts at a particular angle $\theta$ must keep the same angle throughout its path. In Fig. 1, we have found some results for the critical impact parameter from the numerical solution of the orbit Eqs. (7) and (8). In this case, we have obtained the geodesic trajectory (blue thick curve) computed for the impact parameter $b=5.5M$. Therefore, as a direct influence of the existence of the extra-dimensional parameter $\ell$, we observed an increase in the absorption section. Furthermore, in Fig. 1, whereas the black disk depicts the limit of the event horizon, the critical radius for the photon sphere is shown as an internal dotted circle. In turn, the critical impact parameter (black hole shadow) is shown by an external dashed circle.
Figure 1: Null geodesic in polar coordinates for different values of $\ell$, in which we consider a test beam of light by keeping the impact parameter, $b=5.5M$, fixed.
In this scenario, it is important to verify the instability of null circular orbits. For this, we have two important quantities, namely the Lyapunov exponent $\lambda=\sqrt{\mathcal{V}^{\prime\prime}(r)/2\dot{t}^{2}}$ that determines the unstable timescale and the angular velocity $\Omega_{c}=\dot{\phi}/\dot{t}$ for the last null circular orbit (cardoso2009geodesic, ; Anacleto:2021qoe, ; Campos:2021sff, ). The instability exponent is defined by the ratio of the Lyapunov exponent to the angular velocity. In Fig. 2, we see the variation of the instability exponent in relation to the radius $r_{c}$. The maximum of the instability increases with the increase in the contribution coming from the extra dimension. On the other hand, the instability decreases as $r_{c}$ increases. It is worth mentioning that this result is directly sensitive to the geometry, so this analysis is expected to eventually be used to distinguish between Schwarzschild and braneworld black holes.
Figure 2: Lyapunov exponent normalized by the angular velocity $\left(\lambda/\Omega_{c}\right)$, as function of $r_{c}/M$ for some values of $\ell$.
Finally, one can estimate the size of the radius of the black hole shadow by using the study of geodesic curves that we might express via celestial coordinates (vazquez2003strong, )
$\alpha$ $=$ $\lim\limits_{r_{o}\to\infty}\left[-r_{o}^{2}\sin\theta_{o}\dfrac{d\phi}{dr}\Bigr{\rvert}_{\theta=\theta_{o}}\right],$ (11)
$\beta$ $=$ $\lim\limits_{r_{o}\to\infty}\left[r_{o}^{2}\dfrac{d\theta}{dr}\Bigr{\rvert}_{\theta=\theta_{o}}\right],$ (12)
where $\left(r_{o},\theta_{o}\right)$ is the position of the observer at infinity. Thus, for an observer on an equatorial plane $\theta_{o}=\pi/2$, we have the following relationship for the shadow radius
$R_{s}\equiv\sqrt{\alpha^{2}+\beta^{2}}=\sqrt{\dfrac{3r_{c}^{5}}{3r_{c}^{3}-6Mr_{c}^{2}-4M\ell^{2}}}.$ (13)
Let us now expand the radius of the shadow up to order $\left(\ell^{2}/M^{2}\right)$ by assuming that the contribution arising from extra dimension is a small correction to $R_{s}$, i.e., $\ell/M\ll 1$, so that
$R_{s}\approx 3\sqrt{3}M\left[1+\frac{2}{27}\frac{\ell^{2}}{M^{2}}+\mathcal{O}\left(\frac{\ell^{4}}{M^{4}}\right)\right].$ (14)
Notice that the extra-dimensional deviation can be understood as a correction to the Schwarzschild’s shadow radius estimate. In the next section, we will find limits on the correction to the shadow radius.
III Constraints on the $\mathrm{AdS_{5}}$ curvature radius
As we saw above, we obtained deviations in the black hole shadow radius owing to the existence of an extra dimension. On its turn, if we turn off the extra-dimensional effects ($\ell=0$), we recover the well-known results for the Schwarzschild background. Although this independent approach allows us to search for traces of extra dimensions, imposing constraints on deviations found, once the curvature radius $\ell$ is positively correlated to the shadow radius, and given the low-precision measurements, thus we must have a weakening in the predicted bounds compared to sub-mm gravity experiments, for instance.
In our analysis, we aim to obtain bounds on the $\mathrm{AdS_{5}}$ curvature radius. Therefore, the observed angular radius—which has ringlike features in the horizon-scale EHT image of Sgr A*—must be compared to the theoretically determined angular radius of the black hole shadow. From the precise determination of both mass and distance of Sgr A*, the central supermassive black hole of our galaxy, we can verify the accordance, within the uncertainty allowed, between the predicted theoretical results by this study and the empirical data found by the EHT observations (akiyama2022millimeter, ). On its turn, the mass and distance of Sgr A* have been studied over the past few decades by exploring the dynamics of star clusters and, in particular, the movement of stars close to the center of the galaxy. According to Ref. (EventHorizonTelescope:2022xqj, ), the shadow radius measurements are restricted to the following region determined by the estimates based on the Keck and Very Large Telescope Interferometer (VLTI) instruments, within $1\sigma$ confidence interval, respectively,
$4.5\lesssim R_{s}/M\lesssim 5.5,\quad\text{and}\quad 4.3\lesssim R_{s}/M\lesssim 5.3$ (15)
In Fig. 3a, we show the expected deviations in the behavior of the shadow radius owing to the extra dimension. As we see, the increase in the $\mathrm{AdS_{5}}$ curvature radius $\ell$ results in a consequent increase in the radius of the black hole shadow. In our analysis, we have considered the mass from Keck- and VLTI-based estimates (vagnozzi2022horizon, ). Therefore, from the empirical data obtained by observations of EHT, we can impose the upper limits on curvature radius at the $68\%$ confidence level (CL), such that
$\ell_{\left(\text{Keck}\right)}\lesssim 5.7\times 10^{9}\,\mathrm{m}\simeq 3.8\times 10^{-2}\,\mathrm{AU},\quad\text{and}\quad\ell_{\left(\text{VLTI}\right)}\lesssim 2.4\times 10^{9}\,\mathrm{m}\simeq 1.6\times 10^{-2}\,\mathrm{AU},$ (16)
where $1\mathrm{AU}$ is one astronomical unit. These results are shown in Fig. 3b.
At this point, we can compare our results with another which investigates similar physical systems. In this case, the absence of deviations on observation of the highly circular shadow of black hole M87* provides constraints on the $\mathrm{AdS_{5}}$ radius so that $\ell\lesssim 170\,\mathrm{AU}$ (Vagnozzi:2019apd, ). In turn, the first bounds obtained on the extra dimensions by the study of the time lag between the detection of gravitational waves (GW) and electromagnetic (EM) signals from the merging of a binary system detected by the event GW170817 have established at $68\%$ CL an upper limit of $\ell\lesssim 0.535\,\mathrm{Mpc}\simeq 1.10\times 10^{11}\,\mathrm{AU}$ (Visinelli:2017bny, ), whereas our analysis found $\ell_{\left(\text{Keck}\right)}\lesssim 3.8\times 10^{-2}\,\mathrm{AU}\left(\ell_{\left(\text{VLTI}\right)}\lesssim 1.6\times 10^{-2}\,\mathrm{AU}\right)$ for the same confidence interval. It is worth mentioning that, as we have seen, our study yields more stringent constraints than those found by similar physical systems, according to the best of our knowledge. Finally, although we obtained weaker constraints compared to results obtained from sub-mm scale gravity experiments, this approach has set independent limits and can aid the investigation of the geometric structure of the Universe.
Figure 3: (a) Shows the changes in the behavior of the shadow radius with the variation of the $\mathrm{AdS_{5}}$ radius $\ell$. In turn, (b) we found bounds from the direct comparison between the results obtained by our theoretical analysis with the EHT experimental data.
IV Final Remarks
Recently, higher-dimensional theories have drawn increasing attention. As has been pointed out, the seek for evidence of large extra dimensions has been performed in several experimental researches, making its detection a challenging problem in itself. By investigating the null geodesic curves, we find theoretical prediction for the black hole shadow radius, aiming to search for traces of extra dimensions. From recent measurements of the shadow image of Sgr A*—a supermassive black hole in the center of our Galaxy—we estimate deviations owing to an extra dimension in the RSII scenario on the theoretical shadow radius predicted. Given the inaccuracy involving the measured spin of Sgr A*, which may even be zero but still can provide compatible results with recent measures (Fragione:2020khu, ), we explore the possibility of describing them as a nonrotating RSII black hole. In this case, we have considered the Garriga-Tanaka solution that takes a classical approach to linearized braneworld gravity and describes a black hole solution in the weak field limit.
As we saw, the curvature radius of the $\mathrm{AdS_{5}}$ spacetime is positively correlated to the shadow radius so that the increasing/decreasing in one leads to correspondingly increasing/decreasing in the other. Furthermore, we found that $\ell=0$, i.e., in the absence of new physics coming from extra dimension, one recovers the standard estimates expected for the Schwarzschild background. From recent observational data of the shadow of Sgr A* released by the EHT collaboration, one may impose constraints on the curvature radius of $\mathrm{AdS_{5}}$ spacetime. Therefore, by considering the Keck and VLTI estimates, we find more stringent limits on the $\mathrm{AdS_{5}}$ curvature radius, $\ell_{\left(\text{Keck}\right)}\lesssim 3.8\times 10^{-2}\,\mathrm{AU}$ and $\ell_{\left(\text{VLTI}\right)}\lesssim 1.6\times 10^{-2}\,\mathrm{AU}$ (at $68\%$ CL), than previously obtained for similar physical systems, to the best of our knowledge. Besides, a nonzero predicted curvature radius leads to agreement with the observed shadow radius at the $68\%$ confidence interval.
We would like to highlight that, although our analysis provides weaker bounds than those obtained from collider experiments or sub-mm precision gravity tests, which have set limits on the non-Newtonian force, in this study, we have provided conservative bounds since we investigated the shadow radius in a nonrotating RSII-type black hole background. Further, the bounds we reported were obtained independently and in a manner complementary to the previously mentioned studies. Finally, the newly begun era of observational astronomy opens several avenues for future research that might be explored for testing fundamental physics—and so getting new constraints on the $\mathrm{AdS_{5}}$ curvature radius, for instance—or even searching eventual deviations on GR, shedding light into seeking physics beyond the SM.
Acknowledgements.
We thank the referee for the valuable comments and suggestions that have enhanced this work. We would like to thank CNPq, CAPES, and CNPq/PRONEX/FAPESQ-PB (Grant No. $165/2018$), for partial financial support. F.A.B. acknowledges support from CNPq (Grant No. 309092/2022-1). A.S.L. acknowledges support from CAPES (Grant No. 88887.800922/2023-00). J.A.V.C. would like to thank FAPESQ-PB/CNPq (Grant No. 077/2022) for financial support.